m4nt1c0r3:

The radius of a circle is 6cm. The arc length of the sector is 5π cm. What is the measure of the central angle in radians? (write the coefficient as a faction "n/d" without spaces where n is the value of the numerator and d is the value of the denominator. For example: One half would be written as 1/2) What is the area of the sector intercepted by the arc?

snowflake0531:

First we want to find the measure of the central angle in radians The arc length formula is \(s=r(\theta)\) where s is arc length, r is radius, theta is central angle in radians So plugging it in we have \(5\pi=6(\theta)\) So then just solve for theta, but write the answer in the form of a fraction

m4nt1c0r3:

So the central angle would be \[5 \pi/6\] right?

snowflake0531:

Then the area of a sector is \(A=\frac{1}{2}r^2(\theta)\) So plug the theta in from the above, then plug in radius, then find the area

snowflake0531:

@m4nt1c0r3 wrote:
So the central angle would be \[5 \pi/6\] right?
yeah

m4nt1c0r3:

And then the area of the sector would be 16π?

snowflake0531:

\(\frac{1}{2}(6^2)(\frac{5\pi}{6}) = \frac{1}{2}(36)(\frac{5\pi}{6})=18(\frac{5\pi}{6}) = 3(5\pi)=15\pi\)